# From Discrete to Continuous Binary Best-Response Dynamics: Discrete Fluctuations Almost Surely Vanish with Population Size

### v3

@@ VII-J Proof of Theorem 3
(i) Based on 1, the sequence $\langle\bm{x}^{\mathsf{N}}(k)\rangle$ which evolves according to discrete population dynamics (3), is a realization of population dynamics Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ with transition probabilities formulated in (6). (ii) Based on Lemma 1, the collection of $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for the good upper semicontinuous differential inclusion (7). (iii) The transition probabilities of the population dynamics Markov chain (6) are homogeneous and the state space over which the Markov chain is defined is finite for each $\mathsf{N}$. This results in the existence of invariant probability measures ${\mu}^{\frac{1}{\mathsf{N}}}$ for Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ [54]. (iv) 3 specifies the Birkhoff centers of (7). Consider the sequence $\langle\frac{1}{\mathsf{N}}\rangle_{\mathsf{N}\in\mathcal{N}}$ as a vanishing sequence. Theorem 1 and (i)-(iv) together complete the proof. ∎
@@ VII-K Proof of 4
Assume on the contrary that there exists some $N>0$, such that for every population size $\mathsf{N}\in\mathcal{N}_{\geq N}$, where $\mathcal{N}_{\geq N}=\{\mathsf{N}|\mathsf{N}\in\mathcal{N},\mathsf{N}\geq N\}$, the evolution of the proportion of $\mathtt{A}$-players admits some perpetual fluctuation set ${\mathcal{Y}}^{\mathsf{N}}\subset(0,1)\cap\frac{1}{\mathsf{N}}\mathbb{Z}$ of amplitude $y^{\mathsf{N}}$, where $y^{\mathsf{N}}\geq L$ for some positive value $L$,
@@
[…]
where $\mu_{x}^{\frac{1}{\mathsf{N}}}$ is defined as $\mu^{\frac{1}{\mathsf{N}}}\circ\bm{f}^{-1}$, where $\bm{f}:\mathbb{R}^{\mathsf{p}+\mathsf{p}^{\prime}}\to\mathbb{R},$ and $\bm{f}(\bm{x})=\sum_{p=1}^{\mathsf{p}}x_{p}.$ Let $\mathtt{n}$ denote the total number of abstract equilibrium points, and let $q_{p}$, for $p\in[\mathtt{n}]$, denote the $p^{\text{th}}$ equilibrium point of the abstract dynamics. Given Theorem 3, it follows that for every vanishing sequence $\langle\frac{1}{n}\rangle=\frac{1}{n_{0}},\frac{1}{n_{1}},\ldots$ where $n_{i}\in\mathcal{N}$ for $i\in\mathbb{Z}_{\geq 0}$ $\lim_{i\to\infty}\mu_{x}^{\frac{1}{n_{i}}}(\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2))=1,$ for a small enough positive $\epsilon<L/\mathtt{n}$. Equivalently, for every $\delta>0$, there exists $i^{*}>0$ such that for all $i>i^{*}$, $\smash{\mu_{x}^{\frac{1}{n_{i}}}(\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2))>1-\delta}$ and, in turn, $\smash{\mu_{x}^{\frac{1}{n_{i}}}([\min\mathcal{Y}^{\mathsf{N}},\max\mathcal{Y}^{\mathsf{N}}]-\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2))<\delta}$. On the other hand, in view of (20), $[\min\mathcal{Y}^{\mathsf{N}},\max\mathcal{Y}^{\mathsf{N}}]-\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2)$ contains a connected interval of amplitude at least $(L-\mathtt{n}\epsilon)/(\mathtt{n}+1)$, This contradicts (21) for $K=(L-\mathtt{n}\epsilon)/(\mathtt{n}+1)$. ∎
@@
As the update mechanism is asynchronous, the population states at two consecutive time indices $\bm{x}^{\mathsf{N}}(k)$ and $\bm{x}^{\mathsf{N}}(k+1)$ can differ by at most ±$\pm\frac{1}{\mathsf{N}}$. Which element of the population state $\bm{x}^{\mathsf{N}}(k+1)$ may differ from that of $\bm{x}^{\mathsf{N}}(k)$ is determined by the random variable ${P}_{k}$ which is the active subpopulation at time index $k$, and can take values in $[\mathsf{p}+\mathsf{p}^{\prime}]$. The probability of ${P}_{k}=p$ equals $\rho_{p}$. The amount of change in the ${P}_{k}$th element of $\bm{x}^{\mathsf{N}}(k)$ depends on the current strategy of the active agent and her preferred strategy. The former is captured by the random variable ${S}_{k}$ taking values $1$ and $2$, and the latter is determined by the function $s^{*}({P}_{k},\bm{x}^{\mathsf{N}}(k),{S_{k}})$ which is obtained from (1) and (2). The probability of ${S}_{k}=1$ (resp. ${S}_{k}=2$) given the active subpopulation is $p$ is $x^{\mathsf{N}}_{p}/\rho_{p}$ (resp. $1-x^{\mathsf{N}}_{p}/\rho_{p}$). If the current strategy and the preferred strategy are equal, the population states at time indices $k$ and $k+1$ are the same, and if the current strategy is $\mathtt{B}$ and the preferred one is $\mathtt{A}$, the ${P}_{k}$th element of $\bm{x}^{\mathsf{N}}(k)$ increases by $\frac{1}{\mathsf{N}}$, and the ${P}_{k}$th element of $\bm{x}^{\mathsf{N}}(k)$ decreases by $\frac{1}{\mathsf{N}}$ otherwise. This update rule is equivalent to (3). ∎
@@
$\begin{cases}\frac{1}{2}&\text{if }p\leq\mathsf{p},x^{\mathsf{N}}=\tau_{p},\text{ and }s=2,\\ &\text{or }p\leq\mathsf{p},x^{\mathsf{N}}=\tau_{p}+\frac{1}{\mathsf{N}},\text{and }s=1,\\ &\text{if }p>\mathsf{p},{x}^{\mathsf{N}}=\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\text{ and }s=2,\\ &\text{if }p>\mathsf{p},{x}^{\mathsf{N}}=\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}+\frac{1}{\mathsf{N}},\text{ and }s=1,\\ 1&\text{if }p\leq\mathsf{p}\text{ and }x^{\mathsf{N}}<\tau_{p},\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}>\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\text{ and }s=2,\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}>\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}+\frac{1}{\mathsf{N}},\text{ and }s=1,\\ 0&\text{if }p\leq\mathsf{p},x^{\mathsf{N}}>\tau_{p},\text{ and }s=2,\\ &\text{if }p>\mathsf{p}\text{ and }x^{\mathsf{N}}<\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\\ &\text{if }p\leq\mathsf{p},x^{\mathsf{N}}>\tau_{p}+\frac{1}{\mathsf{N}},\text{and }s=1,\end{cases}$
@@
$\begin{cases}\frac{1}{2}&\text{if }p\leq\mathsf{p},x^{\mathsf{N}}=\tau_{p},\text{ and }s=2,\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}=\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\text{ and }s=2,\\ &\text{if }p\leq\mathsf{p},x^{\mathsf{N}}=\tau_{p}+\frac{1}{\mathsf{N}},\text{and }s=1,\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}=\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}+\frac{1}{\mathsf{N}},\text{ and }s=1,\\ 0&\text{if }p\leq\mathsf{p}\text{ and }x^{\mathsf{N}}<\tau_{p},\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}>\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}+\frac{1}{\mathsf{N}},\text{ and }s=1,\\ &\text{if }p>\mathsf{p},x^{\mathsf{N}}>\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\text{ and }s=2,\\ 1&\text{if }p\leq\mathsf{p},x^{\mathsf{N}}>\tau_{p},\text{ and }s=2,\\ &\text{if }p>\mathsf{p}\text{ and }x^{\mathsf{N}}<\tau^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p},\\ &\text{if }p\leq\mathsf{p},x^{\mathsf{N}}>\tau_{p}+\frac{1}{\mathsf{N}},\text{and }s=1.\\ \end{cases}$

### v4

@@ VII-J Proof of Theorem 3
(i) Based on 1, the sequence $\langle\bm{x}^{\mathsf{N}}(k)\rangle$ which evolves according to discrete population dynamics (3), is a realization of population dynamics Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ with transition probabilities formulated in (6). (ii) Based on Lemma 1, the collection of $\langle\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}\rangle_{\mathsf{N}\in\mathcal{N}}$ is a GSAP for the good upper semicontinuous differential inclusion (7). (iii) The transition probabilities of the population dynamics Markov chain (6) are homogeneous and the state space over which the Markov chain is defined is finite for each $\mathsf{N}$. This results in the existence of invariant probability measures ${\mu}^{\frac{1}{\mathsf{N}}}$ for Markov chain $\langle\mathbf{X}^{\frac{1}{\mathsf{N}}}_{k}\rangle_{k}$ [lalley]. (iv) 3 specifies the Birkhoff centers of (7). Consider the sequence $\langle\frac{1}{\mathsf{N}}\rangle_{\mathsf{N}\in\mathcal{N}}$ as a vanishing sequence. Theorem 1 and (i)-(iv) together complete the proof. ∎
@@ VII-K Proof of 4
Assume on the contrary that there exists some $N>0$, such that for every population size $\mathsf{N}\in\mathcal{N}_{\geq N}$, where $\mathcal{N}_{\geq N}\subseteq\{\mathsf{N}|\mathsf{N}\in\mathcal{N},\mathsf{N}\geq N\}$ and unbounded, the evolution of the proportion of $\mathtt{A}$-players admits some perpetual fluctuation set ${\mathcal{Y}}^{\mathsf{N}}\subset(0,1)\cap\frac{1}{\mathsf{N}}\mathbb{Z}$ of amplitude $y^{\mathsf{N}}$, where $y^{\mathsf{N}}\geq L$ for some positive value $L$,
@@
$\text{ and $\mathcal{Z}^{\mathsf{N}}$ connected,}\qquad\mu_{x}^{\frac{1}{\mathsf{N}}}(\mathcal{Z}^{\mathsf{N}})>\delta,$ (22)
where $\mu_{x}^{\frac{1}{\mathsf{N}}}$ is defined as $\mu^{\frac{1}{\mathsf{N}}}\circ f^{-1}$, where $f:\mathbb{R}^{\mathsf{p}+\mathsf{p}^{\prime}}\to\mathbb{R},$ and $f(\bm{x})=\sum_{p=1}^{\mathsf{p}}x_{p}.$ Let $\mathtt{n}$ denote the total number of abstract equilibrium points, and let $q_{p}$, for $p\in[\mathtt{n}]$, denote the $p^{\text{th}}$ equilibrium point of the abstract dynamics. Given Theorem 3, it follows that for every vanishing sequence $\langle\frac{1}{n}\rangle=\frac{1}{n_{0}},\frac{1}{n_{1}},\ldots$ where $n_{i}\in\mathcal{N}$ for $i\in\mathbb{Z}_{\geq 0}$ $\lim_{i\to\infty}\mu_{x}^{\frac{1}{n_{i}}}(\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2))=1,$ for a small enough positive $\epsilon<L/\mathtt{n}$. Equivalently, for every $\delta>0$, there exists $i^{*}>0$ such that for all $i>i^{*}$, $\smash{\mu_{x}^{\frac{1}{n_{i}}}(\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2))>1-\delta}$ and, in turn, $\smash{\mu_{x}^{\frac{1}{n_{i}}}([\min\mathcal{Y}^{\mathsf{N}},\max\mathcal{Y}^{\mathsf{N}}]-\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2))<\delta}$. On the other hand, in view of (21), $\min\mathcal{Y}^{\mathsf{N}},\max\mathcal{Y}^{\mathsf{N}}]-\bigcup_{p=1}^{\mathtt{n}}{\mathcal{B}}({q_{p}},\epsilon/2)$ contains a connected interval of amplitude at least $(L-\mathtt{n}\epsilon)/(\mathtt{n}+1)$, This contradicts (22) for $K=(L-\mathtt{n}\epsilon)/(\mathtt{n}+1)$. ∎
@@
As the update mechanism is asynchronous, the population states at two consecutive time indices $\bm{x}^{\mathsf{N}}(k)$ and $\bm{x}^{\mathsf{N}}(k+1)$ can differ by at most $\pm\frac{1}{\mathsf{N}}$. Which element of the population state $\bm{x}^{\mathsf{N}}(k+1)$ may differ from that of $\bm{x}^{\mathsf{N}}(k)$ is determined by the random variable ${P}_{k}$ which is the active subpopulation at time index $k$, and can take values in $\mathsf{p}+\mathsf{p}^{\prime}]$. The probability of ${P}_{k}=p$ equals $\rho_{p}$. The amount of change in the ${P}_{k}$th element of $\bm{x}^{\mathsf{N}}(k)$ depends on the current strategy of the active agent and her preferred strategy. The former is captured by the random variable ${S}_{k}$ taking values $1$ and $2$, and the latter is determined by the function $s^{*}({P}_{k},\bm{x}^{\mathsf{N}}(k),{S_{k}})$ which is obtained from (1) and (2). The probability of ${S}_{k}=1$ (resp. ${S}_{k}=2$) given the active subpopulation is $p$ is $x^{\mathsf{N}}_{p}/\rho_{p}$ (resp. $1-x^{\mathsf{N}}_{p}/\rho_{p}$). If the current strategy and the preferred strategy are equal, the population states at time indices $k$ and $k+1$ are the same, and if the current strategy is $\mathtt{B}$ and the preferred one is $\mathtt{A}$, the ${P}_{k}$th element of $\bm{x}^{\mathsf{N}}(k)$ increases by $\frac{1}{\mathsf{N}}$, and the ${P}_{k}$th element of $\bm{x}^{\mathsf{N}}(k)$ decreases by $\frac{1}{\mathsf{N}}$ otherwise. This update rule is equivalent to (3). ∎
@@
[…]
@@
[…]